OpenStudy (anonymous):

what is the exact value of cot 5pie/3?

OpenStudy (anonymous):

find the coordinates corresponding to \(\frac{5\pi}{3}\) on the attached cheat sheet the first coordinate is cosine, the second coordinate is sine

OpenStudy (anonymous):

sin(x) / cos(x) = tan(x) sin(5pi/3) = sin(-pi/3) = -sin(pi/3) = -sqrt(3) / 2 cos(5pi/3) = cos(-pi/3) = cos(pi/3) = 1/2 (no, that's not an error on my part) (-sqrt(3) / 2) / (1 / 2) = (-sqrt(3) * 2) / (2 * 1) = -sqrt(3) tan(5pi/3) = -sqrt(3)

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

welcome